OpenStudy (anonymous):

The quotient of a number and 2 is the same as the difference of the number doubled and 3. @Hero so...I just want to check if im right...

OpenStudy (anonymous):

\[N \div 2 = N^2-3\]

OpenStudy (ranga):

I think "number doubled" means 2xN

OpenStudy (anonymous):

>_> a number doubled is a number times it self...

OpenStudy (ranga):

If I double 50 I think it is 100 and not 50 x 50. We can wait for someone else to clarify this matter.

OpenStudy (anonymous):

ahh it is...my bad...Major brain fart...

OpenStudy (ranga):

So N/2 = 2N - 3 N = 4N - 6 6 = 3N N = 2

OpenStudy (anonymous):

\[N \div 2 = N timesN - 3\]

OpenStudy (anonymous):

Pfft...stuipid thing...

OpenStudy (anonymous):

\[N \div 2 = N \times N - 3\]

OpenStudy (ranga):

On the right it should be 2 x N and not N x N

OpenStudy (anonymous):

DAMNIT i did it again!

OpenStudy (anonymous):

\[N \div 2 = N \times 2 - 3\]

OpenStudy (ranga):

Yes. See a few replies above for solution. N = 2

OpenStudy (anonymous):

The larger of two consecutive integers is 7 greater than twice the smaller. Find the integers. Ok...so... for this one i would need to take... wait what?!

hero (hero):

\(\dfrac{n}{2} = 2n - 3\) \(n = 2(2n - 3)\) \(n = 4n - 6\) \(6 = 4n - n\) \(6 = 3n\) \(2 = n\)

hero (hero):

The first number is \(n\) The second (larger) number is \(n + 1\) \(n + 1 = 2n + 7\)

OpenStudy (anonymous):

Gah...idk...ill never pass math D:

hero (hero):

I know it's the "greater than" part that is throwing you off. The mean 7 "more than" rather than using the greater than symbol \(>\)

OpenStudy (anonymous):

*just sobs*

hero (hero):

By the way, in general, if \(n\) is the first number, then \(n + 1\) is the consecutive number.

hero (hero):

or the second number.

hero (hero):

You have to use your own judgment to figure out if they want you to add the seven or use the greater than symbol.

hero (hero):

I'm sorry. Word problems can ruin people's lives.

OpenStudy (anonymous):

*just quite sobbing*

OpenStudy (anonymous):

so...what would the awnser be... *sobs* The larger of two consecutive integers is 7 greater than twice the smaller. Find the integers. 4, 5 -8, -9 -5, -6 -6, -7

hero (hero):

Just solve the equation.

hero (hero):

You need to find \(n\) and \(n + 1\)

OpenStudy (anonymous):

Waaa? ok...so...how do i start...

hero (hero):

Basically you take this equation: \(n + 1 = 2n + 7\) and you solve for \(n\) Whatever result you get, add 1 to it. Then you will have \(n\) and \(n + 1\)

OpenStudy (anonymous):

I dont know how to solve it...its...to...IDK...it doesnt make sense when i look at it...how would i go a bout findning in?

hero (hero):

\(n + 1 = 2n + 7\) \(1 - 7 = 2n - n\) \(-6 = n\) \(-6 + 1 = n + 1\)

OpenStudy (anonymous):

So...>_> -5?

hero (hero):

\(n = -6\) \(n + 1 = -5\)

OpenStudy (anonymous):

*sobs quitely and hugs hero* Thank you D:

hero (hero):

yw